Grade 12Physical Chemistry

3.00 mol of PCl5 kept in 1L closed reaction vessel was allowed to attain equilibrium at 380K. Calculate composition of the mixture at equilibrium. Kc= 1.80 ?

sudhanshu
12 Years agoGrade 12

1 Answer

Prasoon
6 Years ago
Pcl5 =pcl3+cl2
Initial    3.        0.        0
           3-x .       X .       X
 
    Kc= [pcl5][cl2]
                [cl2]
 
 
1.80 = x^2
           3--x
 
Then solve the equation by quadric method and you will get x= 1.59
 
And concentration of pcl5 would be 3-x = 1.49
 
And concentrate of [cl2]  and. [pcl3]  is 1.59
 
Thanks 
 
Thank